OpenStudy (anonymous):

How do you graph y> -2-1/3x?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

it show that in my book. Did I do the problem right?

OpenStudy (anonymous):

should \[x \neq 0\] then one side y=0 x=-1/6 ,,, (-1/6,0) x=1, y=-7/3 ,, (1,-7/3)

OpenStudy (anonymous):

first y=-2-1/3x ok